Circular No. 8760

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2005 YW (LINEAR) An apparently asteroidal object discovered by the LINEAR survey,

and designated 2005 YW (discovery observation tabulated below; cf. MPS 156174, MPEC 2005-Y43, MPO 102263), has been found to show cometary appearance. L. Buzzi, Varese, Italy, reports that 25 co-added 15-s R-band CCD exposures taken with a 0.60-m reflector on Oct. 11.16–11.17 UT show a bright and condensed coma of diameter 15′′ and a wide tail of length ∼ 40′′ in p.a. 300^{o}. E. Guido and G. Sostero report that CCD frames taken remotely on Oct. 11.49–11.50 with a 0.25-m f/3.4 reflector near Mayhill, NM, show a coma of mag 14.7–15.5 with diameter ∼ 15′′ and a broad tail 25′′ long toward p.a. 310^{o}. R. Ligustri, Udine, Italy, reports that his CCD exposures taken in twilight on Oct. 12.2 with a 0.35-m f/5 reflector show a 30′′ coma and a 45′′ tail in p.a. 328^{o}. G. Masi and S. Foglia, Ceccano, Italy, report that CCD frames taken through clouds on Oct. 12.2 with a 0.28-m f/5 reflector show the object to have a 20′′ coma of mag ∼ 15.

2005 UT  α_{2000}  δ_{2000}  Mag.
Dec. 21.27039  5^{h}56^{m}08.^{s}80  +55^{o}10′37′′.0  19.9

Additional astrometry, the following orbital elements, and an ephemeris appear on MPEC 2006-T62:

Epoch = 2006 Dec. 11.0 TT

T = 2006 Dec. 7.8494 TT  ω = 234.^{o}6340
e = 0.989538  Ω = 302.2144 }2000.0
q = 1.993023 AU  i = 40.5438

SUPERNOVA 2006jb IN UGC 12897 Further to IAUC 8758, N. Joubert and W. Li report the LOSS dis-

covery of an apparent supernova on unfiltered KAIT images taken on Oct. 11.25 (at mag 17.4) and 12.29 UT (mag 17.5). The new object is located at α = 0^{h}00^{m}37^{s}.55, δ = +28^{o}23′02′′.9 (equinox 2000.0), which is 4′′.5 west and 1′′.6 south of the nucleus of UGC 12897. A KAIT image taken on Sept. 23.32 showed nothing at this position (limiting mag 19.1).

VY AQUARII CBET 664 contains details of the first reported outburst of VY Aqr in

12 years (at mag ∼ 11 on Oct. 7).

2006 October 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green